OpenStudy (raffle_snaffle):

Find values of t for which vectors a and c are orthogonal. a=<2,-2,4> c=<2t, 0, -t^3>

OpenStudy (raffle_snaffle):

I know we use dot product

OpenStudy (raffle_snaffle):

I already found one of the values, supposedly there is more than one value of t.

OpenStudy (raffle_snaffle):

I get an equation such as 

OpenStudy (anonymous):

If 'a' and 'c' are orthogonal then their dot product is 0. Let's evaluate using the algebraic form of the dot-product and determine the value of 't' accordingly: a · c = 2(2t) + (-2)(0) + (4)(t^3) = 0 => 4t^3 + 4t = 0 Now factor and determine the possible for 't'. @raffle_snaffle

OpenStudy (anonymous):

What do you get for 't'? @raffle_snaffle

OpenStudy (raffle_snaffle):

I used my calculator and got t=0

OpenStudy (raffle_snaffle):

I think my equation is wrong...

OpenStudy (anonymous):

Correct. t = 0 is the right answer. And so it turns out that 'c' is just the zero vector since t = 0. @raffle_snaffle

OpenStudy (raffle_snaffle):

Is there another value of t? My previous professor told me there was another value of t

OpenStudy (anonymous):

There is another value of 't' but it's complex.

OpenStudy (raffle_snaffle):

Well for one thing I didn't get the same equation as you.

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

The equation should -4t^3 + 4t = 0 instead. I forgot the negative.

OpenStudy (raffle_snaffle):

I messed up when you mult. the magnitude of a and c

OpenStudy (anonymous):

Now factoring out -4t we get the following: -4t(t^2 - 1) = 0 But note that t^2 - 1 can further be factored using difference of squares so we get: -4t(t-1)(t+1) = 0 And now we conclude that t = 0 or t = 1 or t = -1. @raffle_snaffle

OpenStudy (raffle_snaffle):